chipper:

Juniors and seniors at a local school were surveyed to see if they were left-handed or right-handed. The data is displayed in the table below: Junior Senior Left-handed 36 14 Right-handed 44 106 If a student is chosen at random from this group, what are the chances of choosing a student who is a left-handed senior? (2 points) Question 2 options: 1) 7% 2) 18% 3) 22% 4) 53%

Narad:

Probability of choosing a left handed senior student is \[=14/170=\] \[=14/170= 0.08\] \[=8%\] The answer is option 1